OpenStudy (ray10):

have no idea how to do this Use the Newton-Raphson method once to find an improved solution to \[f(x) = x \times \cos (x) - \frac{ 1 }{ x }\] . Using an Initial value of \[x _{0} = 4.005^{c} (radians)\]

OpenStudy (anonymous):

https://en.wikipedia.org/wiki/Newton's_method

OpenStudy (ray10):

I believe it is similar to the Approximation technique but I don't know the method on going about it

OpenStudy (anonymous):

Note \(f'(x)=\cos x-x\sin x+\dfrac1{x^2}\)$$x_0=4.005\\x_1=x_0-\frac{f(x_0)}{f'(x_0)}=4.005-\frac{4.005\cos(4.005)-\frac1{4.005}}{\cos(4.005)+4.005\sin(4.005)-\frac1{4.005^2}}\approx3.246$$

OpenStudy (anonymous):

Oops I messed up the signs. $$x_1=4.005-\frac{4.005\cos4.005-\frac1{4.005}}{\cos4.005-4.005\sin4.005+\frac1{4.005^2}}\approx5.166$$

OpenStudy (ray10):

Yes that is what I came to but I could not acquire the approximation. I placed those values into my calculator and it wouldn't come out as a decimal value, instead it came out as I typed it :S

OpenStudy (ray10):

does it lie in the fact that the value is in radians? do I need to change my calculator mode?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=solve+x+cos+x+-+1%2Fx+%3D+0+using+newton%27s+method+starting+at+x_0+%3D+4.005 Click 'Show Details' in the 'Steps' section and you'll see it agrees with me

OpenStudy (anonymous):

Ray10 maybe your calculator is in some 'exact' mode

OpenStudy (ray10):

ah I see, so I had a look and now I tried it again, it came out as what you got too! :) thank you for that! You helped me get my head around that, I was going all over the place with my working out before. One last question; Can this "Newton Method' be seen as the Approximation formula?

OpenStudy (anonymous):

Newton's method is most definitely an iterative approximation algorithm; if you look at the Wolfram Alpha output, you'll see it even shows an 'iteration diagram'.

OpenStudy (anonymous):

It belongs to the larger study known as numerical analysis. https://en.wikipedia.org/wiki/Numerical_analysis

OpenStudy (jhannybean):

Doyou know the calculator method of solving Newton's Method?

OpenStudy (jack1):

@oldrin.bataku mate what did newton use it for...?

OpenStudy (ray10):

I saw that just then @oldrin.bataku , it makes a lot more sense @Jhannybean do you know the way to solve it through calculator?

OpenStudy (anonymous):

Newton used a special case of it to determine the roots of higher-degree polynomials. Here's a good look at how it works: https://upload.wikimedia.org/wikipedia/commons/e/e0/NewtonIteration_Ani.gif

OpenStudy (anonymous):

It uses local information at a point to try and find x-intercepts, so it's known as locally convergent. http://en.wikipedia.org/wiki/Local_convergence

OpenStudy (jack1):

cheers @oldrin.bataku slaters @Jhannybean

OpenStudy (ray10):

@oldrin.bataku thank you a lot for your help!! :) I have to go now, been on a way too long study session! You really helped me out!